Perfect Complements

If 4x,y are perfect complements
then 1x = 4y

If x,3y are perfect complements
then 1x = 3y

Can someone please explain the theory behind this and steps on how to solve related questions like this ?

What would happen if 4x,2y are perfect complements
what would x equal to ?
Answer #1

u Cr, M2) = min (221972) A consumer with this type of utility for only achieves a higher utility level when the quantity of bthen if if u &y are Pe then for solving take 4x = 14 & 8 3y are ac then for solving take 1x = 38. undry are pc then for solviThe following example might be helpful Max u min ( 3 , ²2} 1 Subject to a + b = M. =) 52, + 52 = 60 - 5 (1 +22) = 60 >> Rutny - 12 butting 324 = 22 M + 324 = 12 ->